OpenStudy (anonymous):

please help!? medal rewarded.. What does prefix "milli-" represent? A. 10^-3 B. 10^3 C. 10^4 D. 30

OpenStudy (jamesj):

milli means 1/1000. Hence milligram is one one-thousandth of a gram. Which answer is that?

OpenStudy (anonymous):

C?

OpenStudy (jamesj):

Nope. \[ 10^3 = 1000 \] Hence \[ 10^{-3} = \frac{1}{10^3} = \frac{1}{1000} \]

OpenStudy (anonymous):

lol I was gona say B at first but then had that second thought mindset lol. thanks!

OpenStudy (jamesj):

But B isn't the right answer. A is the correct answer, yes?

OpenStudy (anonymous):

oh really?